P = 737.2228 mmhg

V = 0.164 L

R = 62.4 L*mmHg/ mole*K

T = 21.0oC = 294.0K 

N=?


1
Expert's answer
2021-11-23T16:58:03-0500

Ideal gas law

PV=nRT

"n = \frac{PV}{RT} \\

n = \frac{737.2228 \times 0.164}{62.4 \times 294.0} = 0.0066 \;mol"